# Conditional Denoising Meets Polynomial Modeling: A Flexible Decoupled Framework for Time Series Forecasting

### v5

@@ Abstract
Time series forecasting models are becoming increasingly prevalent due to their critical role in decision-making across various domains. However, most existing approaches represent the coupled temporal patterns, often neglecting the distinction between its specific components. In particular, fluctuating patterns and smooth trends within time series exhibit distinct characteristics. In this work, to model complicated temporal patterns, we propose a Conditional Denoising Polynomial Modeling (CDPM) framework, where probabilistic diffusion models and deterministic linear models are trained end-to-end. Instead of modeling the coupled time series, CDPM decomposes it into trend and seasonal components for modeling them separately. To capture the fluctuating seasonal component, we employ a probabilistic diffusion model based on statistical properties from the historical window. For the smooth trend component, a module is proposed to enhance linear models by incorporating historical dependencies, thereby preserving underlying trends and mitigating noise distortion. Extensive experiments conducted on six benchmarks demonstrate the effectiveness of our framework, highlighting the potential of combining probabilistic and deterministic models. [^2] Our code is open-sourced and available at https://github.com/zjt-gpu/FDF.
@@ 1 Introduction
Time series forecasting is a fundamental task across various domains, playing a crucial role in driving informed decision-making in industries such as finance, healthcare, and manufacturing Kashpruk et al. (2023); Liu et al. (2024). It involves analyzing historical patterns to model the dynamic evolution of underlying processes and predict future trends Montgomery et al. (2015). Accurate time series forecasting is essential for tasks such as demand prediction, resource allocation, enabling organizations to anticipate future needs and optimize operations. Trastour et al. (2002); Kilcioglu et al. (2017); Drutsa et al. (2015)
@@
Figure 1: The forward diffusion process iteratively adds noise to the original time series, which causes the trend components to be overshadowed as the noise increasingly dominates. This makes it challenging to recover trend during the reverse denoising process.
@@
Over the years, various approaches have been proposed to address this challenge, ranging from classical statistical models, such as Autoregressive Integrated Moving Average (ARIMA) Asteriou and Hall (2011), to machine learning techniques Masini et al. (2023), and advanced deep learning models, including recurrent neural networks (RNNs) Lin et al. (2023); Lai et al. (2018), convolutional neural networks (CNNs) Bai et al. (2018); Koprinska et al. (2018), and transformer-based models Zhou et al. (2021); Wu et al. (2021). Large Language Models (LLM) like Time-LLM and LLMTIME Jin et al. (2023); Gruver et al. (2024) have demonstrated the potential in time series forecasting. In addition to their success, probabilistic models have emerged as promising alternatives, offering a principled framework for quantifying uncertainty and modeling complex distributions. Diffusion models Rasul et al. (2021); Shen and Kwok (2023), becoming the leading paradigm for probabilistic models originally developed for generation, have proven effective in time series forecasting due to their ability to capture intricate stochastic processes.
@@
Diffusion-based methods have shown significant potential in time series forecasting. We argue that these approaches are critically limited by the naive application of noise addition and denoising processes, which fail to account for the distinct dynamics of different patterns. Specifically, noise is uniformly added across the entire time series, and denoising is then applied indiscriminately, treating all temporal components equally. However, time series inherently exhibits diverse dynamics, including trends and seasonal patterns, which require distinct treatment. Trends patterns, reflecting gradual shifts over extended periods, and seasonal patterns, characterized by periodic fluctuations, possess different structural properties. A uniform noise addition and denoising strategy fails to account for these differences, potentially distorting key components, particularly intrinsic trends and fluctuations, which are crucial for accurate forecasting. As shown in Fig.1, this indiscriminate approach disrupts critical trend information during denoising, hindering trend recovery and reducing forecasting accuracy.
@@
Employing suitable models to capture the different components is crucial. Single model often struggles to capture the subtle and dynamic interplay between trends and variations, which can significantly affect the model’s forecasting ability. Diffusion models, with their advantages in modeling complex dynamic processes, capturing high-frequency fluctuations, and quantifying uncertainty, are more suitable for handling components characterized by large fluctuations and high-frequency variations in time series. In contrast, deterministic models, such as models based on linear layers, learn features from historical patterns and are particularly effective at identifying regular patterns, making them better suited for modeling components with stable trends. Therefore, by leveraging the strengths of both model types and tailoring them to the distinct characteristics of time series, we can more comprehensively capture the dynamic variations within the series and enhance forecasting accuracy.
@@
To overcome these challenges, we propose a novel Conditional Denoising Polynomial Modeling (CDPM) framework that leverages the inductive bias of time series decomposition. By separating the time series into trend and seasonal components, we enable decoupled modeling tailored to each component’s characteristics. Specifically, we employ carefully designed Conditional Denoising Seasonal Module (CDSM) to capture the high-frequency seasonal component, introducing an innovative denoising strategy conditioned on statistical properties derived from historical window. Concurrently, we model the intrinsic trend component using Polynomial Trend Module (PTM), which are adept at capturing gradual changes over time. Additionally, we reformulate the Improved Evidence Lower Bound (IELBO) to facilitate the joint training of the CDSM and PTM. This decoupling method enhances the framework’s ability to capture temporal dependencies and complex patterns. The key contributions of this work are:
@@ 2.1 Deterministic Forecasting
Deterministic models map deterministic inputs to deterministic outputs by learning consistent patterns within time series, making them crucial for capturing predictable temporal relationships. Recurrent neural networks (RNNs), such as LSTNet Lai et al. (2018) and SegRNN Lin et al. (2023), became popular for sequential dependencies, while convolutional neural networks (CNNs) like TCN Bai et al. (2018) and SCINet Liu et al. (2022) excelled at capturing both local and global temporal features. Graph-based models, such as MTGNN Wu et al. (2020) and FourierGNN Yi et al. (2024), improved forecasting by modeling interdependencies between time series. More recently, Transformer-based models like Informer Zhou et al. (2021) and Autoformer Wu et al. (2021) have utilized self-attention mechanisms to efficiently capture long-range dependencies, excelling in high-dimensional and long-horizon tasks. Surprisingly, simpler models like DLinear Zeng et al. (2023) and TSMixer Ekambaram et al. (2023) have shown that well-tuned linear layers, focusing on long-term trends, can outperform complex frameworks, highlighting their efficiency. Despite advances, deterministic models often struggle with volatile fluctuations, limiting their ability to model complex time series dynamics.
2.2 Probabilistic Forecasting
@@
Probabilistic models offer uncertainty estimation and the ability to model temporal distributions. These models, including Variational Autoencoders (VAEs)Kingma and Welling (2022), Generative Adversarial Networks (GANs)Goodfellow et al. (2014), and flow-based models Rezende and Mohamed (2016), are designed to capture high-dimensional distribution. Recently, Denoising Diffusion Probabilistic Models (DDPMs) have emerged as a promising paradigm for generative modeling, demonstrating their potential in various time series applications. DDPMs have advanced time series forecasting by capturing complex temporal patterns. Early models like TimeGrad Rasul et al. (2021) introduced autoregressive denoising with Langevin sampling for multivariate predictions. TSDiff Kollovieh et al. (2024) improved short-term accuracy through self-guiding mechanisms, while score-based models such as ScoreGrad Yan et al. (2021) applied stochastic differential equations (SDEs) for continuous-time forecasting, expanding diffusion to fluctuatingly sampled data. Conditional models like TimeDiff Shen and Kwok (2023) incorporated external information to guide the diffusion process, outperforming traditional methods. Latent Diffusion Models (LDMs)Rombach et al. (2022) enhanced efficiency by performing diffusion in lower-dimensional spaces. This approach was further demonstrated by LDCast and Latent Diffusion Transformers (LDT)Feng et al. (2024), which improved precipitation and time series forecasting. Models like DSPD and CSPD Bilos et al. (2023) applied diffusion to anomaly detection and interpolation. Diffusion models have also been successfully used in flood forecasting Shao et al. (2024), stock market prediction Daiya et al. (2024), and EV load forecasting Li et al. (2024).
@@ 4.2 Preprocessing for CDPM
Instance normalization across both historical and predicted windows is essential for maintaining the stability of the CDPM. We normalize the historical window by subtracting the mean and dividing by the standard deviation Kim et al. (2021), and transfer the same statistics to normalize the predicted window during training. Specifically, $x_{L}$ is normalized to $x_{L}^{0}$, and $x_{T}$ is normalized to $x_{T}^{0}$ using the statistics derived from $x_{L}$. This ensures alignment between the windows and preserves both the temporal structure and stability in the forecasting process. The subsequent denormalization step restores the forecasted values to their original distribution, thereby maintaining consistency with input data.
@@ 4.3 Conditional Denoising Seasonal Module
Modeling fluctuating seasonal components in time series is challenging due to their non-stationary and intricate patterns. We propose the CDSM, based on denoising diffusion model, which is a popular paradigm in probabilistic modeling. By incorporating historical statistical properties as conditional inputs, CDSM captures seasonal patterns more effectively.
@@
Reversed Conditional Denoise.
@@
$\text{AdaLN}(\mathbf{h})=\alpha_{t}\cdot\text{LayerNorm}(\mathbf{h})+\beta_{t}.$ (5)
@@
where $\text{MLP}_{\mu}(\cdot)$ and $\text{MLP}_{\sigma}(\cdot)$ capture the mean and variance of historical window features, respectively. The predicted statistics are used in the denoising process via Gaussian sampling:
@@
$X_{\theta}\left(X_{T,s}^{k},t\mid\nabla_{\mathbf{x}_{T,s}}^{k-1}\right)=\rho_{1}\hat{x}_{T,s}^{0}+\rho_{2}\nabla_{\mathbf{x}_{T,s}}^{k-1}.$ (11)
@@ 4.5 Reconstruction and Optimization
The final prediction $\hat{x}_{T}^{0}$ calculated by integrating the trend and seasonal components as follows:
@@
where $X_{\theta}(\cdot)$ represents the CDSM and $f_{\theta}(\cdot)$ corresponds to the PTM.
Improved Evidence Lower Bound Optimization.
The final prediction $\hat{x}_{T}^{0}$ is subsequently denormalized to $\hat{x}_{T}$ using the mean and variance derived from the historical window, ensuring that the forecast is consistent with the original distribution of the time series. The loss function in CDPM is derived from the Evidence Lower Bound (ELBO). Furthermore, we extend the traditional ELBO to accommodate the separate modeling of the trend and seasonal components. This decoupled modeling of the trend and seasonal components ensures that both contribute independently to the final forecast. Through derivation, we find that by modeling the seasonal component with CDSM and the trend component with PTM, the results can be jointly optimized and trained to improve the overall performance of our framework, leading to more accurate predictions.
The Improved Evidence Lower Bound is given by:
@@ 5.2 Experiment Results
The experimental results in Table 2 strongly validate the effectiveness of the CDPM for time series forecasting. CDPM’s decoupling strategy, which separates trend and seasonal components, allows for more accurate modeling of temporal patterns. It performs particularly well on challenging datasets like ETTh1 and ETTh2, where traditional models such as Diffusion-TS and D3VAE struggle with varying time resolutions and high noise. CDPM’s strength lies in its Polynomial Trend Module (PTM) and Conditional Denoising Seasonal Module (CDSM), which capture complex trends and seasonal dynamics. On high-dimensional datasets like Electricity and Weather, CDPM maintains superior accuracy. Unlike CSDI and Diffusion-TS, which struggle with long forecast horizons, CDPM handles long-term dependencies effectively. On volatile datasets like Wind, CDPM captures sharp fluctuations and achieves competitive results, though further refinement could improve performance on highly noisy data. Overall, CDPM shows strong generalization across various datasets, outperforming models like DLinear and Autoformer. In summary, CDPM’s decoupling strategy enhances performance, confirming the theoretical hypothesis that it improves accuracy and interpretability over traditional and diffusion-based models.
@@
  • w/o CR: Removes the conditional information.
  • PLR: Replaces the PTM with a simple linear layer.
  • w/o INA: Removes instance normalization alignment.
  • w/o CR+PLR: Removes the conditional information and replaces PTM with a linear layer.
@@
Figure 3: Parameter Sensitivity Analysis of CDPM: Hidden Dimension (Left), Patch Length (Middle), and Predicted Series Length (Right).
@@
Decoupling Validation Study.
@@
w/o DECOUPLING 0.3839 0.4106 0.2518 0.3157
REV 0.3541 0.3832 0.2452 0.3122
  • w/o DECOUPLING: This variant directly applies the CDSM to the original time series.
  • REV: In this variant, PTM and CDSM are reversed, with PTM predicting the seasonal component and CDSM predicting the trend component.
The decoupling strategy, which separates trend and seasonal patterns, significantly enhances performance, with CDPM outperforming all ablated models. The improvement is most pronounced on ETTh2, where clear trend-seasonal separation aids effective modeling. On ETTm2, the gain is more modest due to finer granularity and overlapping patterns, but CDPM still outperforms the ablated models, demonstrating its adaptability. Applying the CDSM directly underperforms due to its inability to separate trend and seasonal components, while reversing PTM and CDSM shows slight improvements but still falls short of CDPM, highlighting the importance of proper module assignment. These results emphasize the need to tailor modules to capture distinct trend and seasonal patterns, with decoupling especially beneficial for datasets with clear temporal structures. For finer-grained datasets like ETTm2, the gains from decoupling are more modest, but CDPM consistently outperforms the ablated models, showcasing its robustness across datasets with varying granularities.
@@
To compare the performance of the specially designed components, PTM and CDSM, in modeling the trend and seasonal components of the ETTh1 dataset, we conducted experiments with an input length of 192 and a prediction horizon of 168. The results highlight the effectiveness of tailored modeling strategies for each component and the benefits of decoupled modeling in improving forecasting accuracy. As shown in Figure 4 (a), PTM excels in capturing intrinsic and gradual trends due to its linear structure, accurately modeling patterns and tracking the overall direction of the time series. In contrast, applying CDSM to the trend component introduces unnecessary noise and overfitting to minor variations, diminishing trend quality. For the seasonal component, characterized by complex fluctuations, CDSM outperforms PTM, as shown in Figure 4(b). CDSM effectively captures sharp seasonal changes, while PTM focuses on underlying regularities in the fluctuations, though it may sometimes miss rapid shifts. These results validate each module’s modeling capabilities for its designated component. By assigning each module to the component it best models, CDPM captures both distinct trends and complex fluctuations, enhancing forecasting accuracy, as seen in both trend and seasonal visualizations.
Hyper-parameter Sensitiviy Analysis.
To investigate the impact of key design choices in CDPM, such as hidden dimension size ($d_{\text{model}}$), patch length ($H$), and prediction horizon, on model performance, we performed a hyperparameter sensitivity analysis using the ETTh1 dataset. The results show that increasing $d_{\text{model}}$ initially improves performance, peaking at $d_{\text{model}}=256$, but larger values (e.g., $d_{\text{model}}=512$) yield diminishing returns and potential overfitting, highlighting the trade-off between model capacity and efficiency. Similarly, $H$ significantly affects performance, with $H=8$ capturing both local dependencies and global temporal patterns, while larger values introduce redundancy and reduce efficiency. Finally, CDPM performs consistently across prediction horizons, with PTM capturing trends and CDSM adapting to fluctuations, ensuring reliable forecasting. Overall, the analysis confirms CDPM’s balance between complexity and scalability, making it suitable for diverse forecasting tasks.
@@
To evaluate the framework’s interpretability and modeling capability, we performed a combined performance analysis of the PTM and CDSM within the CDPM, as shown in Figure 5. The CDPM decouples multidimensional time series, capturing diverse trends and cyclic patterns, demonstrating its ability to model complex temporal dynamics. In Figure 5 (left), applied to the Electricity dataset with distinct periodic and trend characteristics, PTM captures intrinsic trends, with predictions closely matching actual trends. Figure 5 (right) highlights CDSM’s ability to model seasonal variations, with predicted components aligning well with actual patterns, compensating for PTM’s trend deviations and improving accuracy. The joint modeling of PTM and CDSM within the CDPM enhances both interpretability and predictive accuracy for datasets with trend and seasonal components.

### v6

@@ Abstract
Time series forecasting models are becoming increasingly prevalent due to their critical role in decision-making across various domains. However, most existing approaches represent the coupled temporal patterns, often neglecting the distinction between their specific components. In particular, fluctuating patterns and smooth trends within time series exhibit distinct characteristics. In this work, to model complicated temporal patterns, we propose a Conditional Denoising Polynomial Modeling (CDPM) framework, where probabilistic diffusion models and deterministic linear models are trained end-to-end. Instead of modeling the coupled time series, CDPM decomposes it into trend and seasonal components for modeling them separately. To capture the fluctuating seasonal component, we employ a probabilistic diffusion model based on statistical properties from the historical window. For the smooth trend component, a module is proposed to enhance linear models by incorporating historical dependencies, thereby preserving underlying trends and mitigating noise distortion. Extensive experiments conducted on six benchmarks demonstrate the effectiveness of our framework, highlighting the potential of combining probabilistic and deterministic models. [^2] Our code is available at https://github.com/zjt-gpu/CDPM.
@@ 1 Introduction
Time series forecasting is a fundamental task across various domains, playing a crucial role in driving informed decision-making in industries such as finance, healthcare, and manufacturing Kashpruk et al. (2023). It involves analyzing historical patterns to model the dynamic evolution of underlying processes and predict future trends Cheng et al. (2025a). Accurate time series forecasting is essential for tasks such as demand prediction, resource allocation, enabling organizations to anticipate future needs and optimize operations Kilcioglu et al. (2017); Liu et al. (2024).
@@
Figure 1: The forward diffusion process iteratively adds noise to the original time series, which causes the trend components to be overshadowed as the noise increasingly dominates.
@@
Over the years, various approaches have been proposed to address this challenge Cheng et al. (2025b), ranging from classical statistical models, such as Autoregressive Integrated Moving Average (ARIMA) Asteriou and Hall (2011), to machine learning techniques Masini et al. (2023), and advanced deep learning models, including recurrent neural networks (RNNs) Lin et al. (2023); Lai et al. (2018), convolutional neural networks (CNNs) Bai et al. (2018); Koprinska et al. (2018), and transformer-based models Zhou et al. (2021); Wu et al. (2021). Large language models (LLMs) such as TimeLLM and LLMTIME Jin et al. (2023); Gruver et al. (2024) have demonstrated the potential in time series forecasting. In addition to their success, probabilistic models have emerged as promising alternatives, offering a principled framework for quantifying uncertainty and modeling complex distributions. Diffusion models Rasul et al. (2021); Shen and Kwok (2023), becoming the leading paradigm for probabilistic models originally developed for generation, have proven effective in time series forecasting due to their ability to capture intricate stochastic processes.
@@
Diffusion-based methods Lee et al. (2024) have shown significant potential in time series forecasting. We argue that these approaches are critically limited by the naive application of noise addition and denoising processes, which fails to account for the distinct dynamics of different patterns. Specifically, noise is uniformly added across the entire time series, and denoising is then applied indiscriminately, treating all temporal components equally. However, time series inherently exhibits diverse dynamics, including trends and seasonal patterns, which require different treatment. Trends patterns, reflecting gradual shifts over extended periods, and seasonal patterns, characterized by periodic fluctuations, possess different structural properties. This strategy of uniform noise addition and denoising fails to account for these differences, potentially distorting key components, particularly intrinsic trends and fluctuations, which are crucial for accurate forecasting. As shown in Fig.1, this indiscriminate approach disrupts critical trend information during denoising, hindering trend recovery and reducing accuracy.
@@
Typically, a single model often struggles to capture the subtle and dynamic interplay between trends and variations, which can significantly affect the model’s overall forecasting ability. Diffusion models, which excel at modeling complex dynamic processes, capturing high-frequency fluctuations, and quantifying uncertainty, are more suitable for handling components characterized by large fluctuations and high-frequency variations in time series. In contrast, deterministic models, such as models based on linear layers, learn features from historical patterns and are particularly effective at identifying regular patterns, making them better suited for modeling components with stable trends. Therefore, by strategically leveraging the strengths of both model types and tailoring them to the distinct characteristics of time series, we can more fully capture the dynamic variations and enhance the final forecasting results.
@@
To overcome these challenges, we propose a novel Conditional Denoising Polynomial Modeling (CDPM) framework that leverages the inductive bias of time series decomposition. By separating the time series into trend and seasonal components, we enable decoupled modeling tailored to the characteristics of each component. Specifically, we employ a carefully designed Conditional Denoising Seasonal Module (CDSM) to capture the high-frequency seasonal component, introducing an innovative denoising strategy conditioned on statistical properties derived from historical window. Currently, we model the intrinsic trend component using the polynomial trend module (PTM), which is adept at capturing gradual changes. Additionally, we reformulate the Improved Evidence Lower Bound (ELBO) to facilitate joint training of the CDSM and PTM. This decoupling method of our framework improves the ability to capture temporal dependencies and complex patterns. The key contributions of this work are:
@@ 2.1 Deterministic Forecasting
Deterministic models map deterministic inputs to deterministic outputs by learning consistent patterns within time series, making them crucial for capturing predictable temporal relationships. Recurrent neural networks (RNNs), such as LSTNet Lai et al. (2018) and SegRNN Lin et al. (2023), became popular for sequential dependencies, while convolutional neural networks (CNNs) Cheng et al. (2024) like TCN Bai et al. (2018) and SCINet Liu et al. (2022) excelled at capturing both local and global temporal features. Graph-based models, such as MTGNN Wu et al. (2020) and FourierGNN Yi et al. (2024), improved forecasting by modeling interdependencies between time series. More recently, Transformer-based models like Informer Zhou et al. (2021) and Autoformer Wu et al. (2021) have utilized self-attention mechanisms to efficiently capture long-range dependencies, excelling in high-dimensional and long-horizon tasks. Surprisingly, well-tuned linear models such as DLinear Zeng et al. (2023) and TSMixer Ekambaram et al. (2023) have demonstrated strong performance by effectively capturing long-term trends, often outperforming more complex architectures and showcasing their efficiency. Despite advances, deterministic models often struggle with volatile fluctuations, limiting their ability to model complex time series dynamics.
2.2 Diffusion-Based Time Series Forecasting
@@
Probabilistic models offer uncertainty estimation and the ability to model temporal distributions. These models, including Variational Autoencoders (VAEs)Kingma and Welling (2022), Generative Adversarial Networks (GANs)Goodfellow et al. (2014), and flow-based models Rezende and Mohamed (2016), are designed to capture high-dimensional distribution. Recently, Denoising Diffusion Probabilistic Models (DDPMs) have emerged as a promising paradigm for generative modeling, demonstrating their potential in various time series applications. DDPMs have advanced time series forecasting by capturing complex temporal patterns. Early models like TimeGrad Rasul et al. (2021) introduced autoregressive denoising with Langevin sampling for multivariate predictions. TSDiff Kollovieh et al. (2024) improved short-term accuracy through self-guiding mechanisms, while score-based models such as ScoreGrad Yan et al. (2021) applied stochastic differential equations (SDEs) for continuous-time forecasting, expanding diffusion to fluctuatingly sampled data. Conditional models like TimeDiff Shen and Kwok (2023) incorporated external information to guide the diffusion process, outperforming traditional methods. Latent Diffusion Models (LDMs)Rombach et al. (2022) enhanced efficiency by performing diffusion in lower-dimensional spaces. This approach was further demonstrated by LDCast and Latent Diffusion Transformers (LDT)Feng et al. (2024), which improved precipitation and time series forecasting. Models like DSPD and CSPD Bilos et al. (2023) applied diffusion to anomaly detection and interpolation.
@@ 4.2 Temporal Feature Structuring
Instance normalization across both historical and predicted windows is essential for maintaining the stability of the CDPM. We normalize the historical window by subtracting the mean and dividing by the standard deviation Kim et al. (2021), and transfer the same statistics to normalize the predicted window during training. Specifically, $x_{L}$ is normalized to $x_{L}^{0}$, and $x_{T}$ is normalized to $x_{T}^{0}$ using the statistics derived from $x_{L}$. This ensures alignment between the windows and preserves both the temporal structure and stability in the forecasting process. The subsequent denormalization step restores the predicted values to their original scale, ensuring consistency with the input data distribution.
@@ 4.3 Conditional Denoising Seasonal Module
Modeling fluctuating seasonal components in time series is challenging due to their non-stationarity and intricate patterns. We propose CDSM, based on a denoising diffusion model, a popular paradigm in probabilistic modeling. By incorporating historical statistical properties as conditional inputs, CDSM captures seasonal patterns more effectively.
@@
Reverse Conditional Denoising Process.
@@
$\text{AdaLN}(\mathbf{h})=\alpha_{t}\cdot\text{LayerNorm}(\mathbf{h})+\beta_{t},$ (5)
@@
where $\text{MLP}_{\mu}(\cdot)$ and $\text{MLP}_{\sigma}(\cdot)$ capture the mean and variance of the historical window features, respectively. These predicted statistics are subsequently used in the denoising process via Gaussian sampling:
@@
$X_{\theta}\left(X_{T,s}^{k},t\mid\nabla_{\mathbf{x}_{T,s}}^{k-1}\right)=\rho_{1}\hat{x}_{T,s}^{0}+\rho_{2}\nabla_{\mathbf{x}_{T,s}}^{k-1},$ (11)
@@ 4.5 Optimization
The final prediction $\hat{x}_{T}^{0}$ is calculated by integrating the trend and seasonal components as follows:
@@
where $X_{\theta}(\cdot)$ represents the CDSM and $f_{\theta}(\cdot)$ corresponds to the PTM, respectively.
Improved ELBO Optimization.
The final prediction $\hat{x}_{T}^{0}$ is subsequently denormalized to $\hat{x}_{T}$ using the mean and variance derived from the historical window, ensuring that the forecast is consistent with the original distribution of the time series. The loss function in CDPM is derived from the Evidence Lower Bound (ELBO). Furthermore, we extend the traditional ELBO to accommodate the separate modeling of the trend and seasonal components. This decoupled modeling of the trend and seasonal components ensures that both contribute independently to the final forecast. Through derivation, we find that by modeling the seasonal component with CDSM and the trend component with PTM, the results can be jointly optimized and trained to improve the overall performance of our framework, leading to more accurate predictions. The improved ELBO is given by:
@@ 5.2 Experiment Results
The experimental results in Table 2 validate the effectiveness of the CDPM for time series forecasting. CDPM’s decoupling strategy, which separates trend and seasonal components, allows for more accurate modeling of temporal patterns. It performs well on challenging datasets like ETTh1 and ETTh2, where traditional models such as Diffusion-TS and D3VAE struggle with varying time resolutions and high noise. CDPM’s strength lies in its Polynomial Trend Module (PTM) and Conditional Denoising Seasonal Module (CDSM), which capture complex trends and seasonal dynamics. On high-dimensional datasets like Electricity and Weather, CDPM maintains superior accuracy. Unlike CSDI and Diffusion-TS, which struggle with long forecast horizons, CDPM handles long-term dependencies effectively. On volatile datasets like Wind, CDPM captures sharp fluctuations and achieves competitive results, though further refinement could improve performance on highly noisy data. Overall, CDPM shows strong generalization across various datasets. In summary, CDPM’s decoupling strategy enhances performance, confirming the theoretical hypothesis that it improves accuracy and interpretability over traditional and diffusion-based models.
@@
  • CDPM1: Removes the conditional information.
  • CDPM2: Replaces the PTM with a simple linear layer.
  • CDPM3: Removes instance normalization alignment.
  • CDPM4: Removes the conditional information and replaces PTM with a linear layer.
@@
Figure 3: Parameter sensitivity analysis of CDPM: hidden dimension (left), patch length (middle), and predicted length (right).
@@
Evaluation of the Decoupling Strategy.
  • CDPM5: This variant directly applies the CDSM to the original time series.
  • CDPM6: In this variant, PTM and CDSM are reversed, with PTM predicting the seasonal component and CDSM predicting the trend component.
The decoupling strategy, which separates trend and seasonal patterns, significantly enhances performance, with our framework CDPM outperforming all ablated models. The improvement is most pronounced on ETTh2, where clear trend-seasonal separation aids effective modeling. On ETTm2, the gain is more modest due to finer granularity and overlapping patterns, but CDPM still outperforms the ablated models, demonstrating its adaptability. Applying our framework directly underperforms due to its inability to separate trend and seasonal components, while reversing PTM and CDSM shows slight improvements but still falls short of our framework, highlighting the importance of proper module assignment. These results emphasize the need to tailor modules to capture distinct trend and seasonal patterns, with decoupling especially beneficial for datasets with clear temporal structures. For finer-grained datasets like ETTm2, the gains from decoupling are more modest, but our framework consistently outperforms the ablated models, showcasing its robustness across datasets with varying granularities.
@@
CDPM5 0.3839 0.4106 0.2518 0.3157
CDPM6 0.3541 0.3832 0.2452 0.3122
@@
To compare the performance of the specially designed components, PTM and CDSM, in modeling the trend and seasonal components of the ETTh1 dataset, we conducted experiments with an input length of 192 and a prediction horizon of 168. The results highlight the effectiveness of tailored modeling strategies for each component and the benefits of decoupled modeling in improving forecasting accuracy. As shown in Figure 4 (a), PTM excels in capturing intrinsic and gradual trends due to its linear structure, accurately modeling patterns and tracking the overall direction of the time series. In contrast, applying CDSM to the trend component introduces unnecessary noise and overfitting to minor variations, diminishing trend quality. For the seasonal component, which exhibits complex fluctuations, the seasonal module demonstrates superior performance, as shown in Figure 4 (b). CDSM effectively captures sharp seasonal changes, while PTM focuses on underlying regularities in the fluctuations, though it may sometimes miss rapid shifts. These results validate each module’s modeling capabilities for its designated component.
Hyper-Parameter Sensitivity Analysis.
To investigate the impact of key design choices in CDPM, such as hidden dimension size ($d_{\text{model}}$), patch length ($H$), and prediction horizon on model performance, we conducted a hyperparameter sensitivity analysis using the ETTh1 dataset. The results show that increasing $d_{\text{model}}$ initially improves performance, peaking at $d_{\text{model}}=256$, while larger values (e.g., $d_{\text{model}}=512$) offer diminishing returns and risk overfitting, highlighting the trade-off between model capacity and efficiency. Similarly, $H$ significantly affects performance, with $H=8$ capturing both local dependencies and global patterns, whereas larger values introduce redundancy and reduce efficiency. Finally, our framework performs consistently across prediction horizons, with the trend module capturing long-term structure and the seasonal module adapting to fluctuations, ensuring reliable forecasts. Overall, the analysis confirms the balance between complexity and scalability in our framework, making it suitable for diverse forecasting tasks.
@@
To assess the interpretability and predictive performance of the framework, we conducted a joint analysis of the PTM and CDSM modules on the Electricity dataset, as shown in Figure 5. The framework decouples multidimensional time series, capturing diverse trends and cyclic patterns, demonstrating its capacity to model complex temporal dynamics. In Figure 5 (left), applied to the Electricity dataset with clear periodic and trend structures, the trend module captures the underlying patterns, with predictions closely matching actual values. Figure 5 (right) illustrates the seasonal module’s ability to model periodic variations, where predicted components align well with the observed sequences, compensating for deviations in trend estimation and improving accuracy. The integration of deterministic and diffusion-based modules within the framework enhances both interpretability and predictive accuracy across datasets with combined components.